### v1

@@ V-A2 Global Agent
The global agent coordinates with pre-trained local agents to achieve system-wide optimization by considering global objectives. Its features include:
  • Aggregate Observability: The global agent has an abstract view of the platform provided by local agents, that we detail in Sect.V-B.
  • Joint Training: The global agent is trained jointly with local agents to align local and global objectives, ensuring both neighborhood-level efficiency and global consistency.
  • Non-Local Reward The non-local reward function, $R_{non-local}$, may include terms such as:
[…]
  • Global Reward The global agent optimizes a reward function that considers both local and non-local scope:
@@
    Where:
    • $R_{non-local}$: Global objective based on system-wide metrics (e.g. overall resource utilization, global SLA adherence, and minimized network congestion).
    • $R_{local_{i}}$: Local rewards from individual local agents.
    • $\lambda,\mu_{i}$: Weight parameters for combining global and local objectives (deciding between linear or nonlinear combinations is part of the design choice).
In future work, we might consider the integration of local and global rewards can be achieved through either linear or nonlinear combinations. The choice of combination strategy will depend on the specific application requirements, such as the need for rapid local responsiveness versus long-term global optimization. This hybrid approach ensures scalability while maintaining overall system efficiency.
@@ V-B Observation Space
The observation space for agents in the Cloud-Edge Continuum consists of two key components: the state of the application graph and the state of the resource graph. Each local agent has partial observability, limited to its assigned local resource network, while the global agent aggregates information from all local agents.
V-B1 State of the Application Graph
The state of the application graph observed by an agent is defined as:
@@ V-B1 State of the Application Graph
V-B2 State of the Resource Graph
For the Cloud-Edge Continuum system, partitioned by geographical locality or different resource providers, each local agent observes only a partial resource graph corresponding to its neighborhood:
@@ V-B2 State of the Resource Graph
where $G_{R}^{local}$ is the subgraph of $G_{R}$ representing the local agent’s assigned resource network and $v_{avail_{i}}=\{cpu(v_{i}),gpu(v_{i}),ram(v_{i}),stor(v_{i}),pt_{i},aval_{i}\}$ expresses the resource availability status. We denote local environment $i$ with $S_{res}^{local}(i)$.
V-B3 Combined Observation Space for Local Agents
The combined observation space for a local agent includes the application graph state and the partial resource graph state:
$O_{local}=\{S_{app},S_{res}^{local}\}$ (11)
V-B4 Combined Observation Space for the Global Agent
The global agent observes aggregated information from all local agents, representing the aggregated full resource graph and application graph:
@@ V-B4 Combined Observation Space for the Global Agent
Partial observation by local agents promotes decentralization, reducing computational complexity by limiting their focus to local subgraphs. This localized approach enables faster, more adaptive decision-making tailored to the immediate environment. Meanwhile, the global agent utilizes aggregated, more abstract representations of local subgraph information, which streamlines its decision-making process. This approach allows the global agent to focus on high-level coordination and system-wide optimization while leveraging the decentralized operations of local agents for detailed, localized management.
@@ V-C Action Space
V-C1 Action Space for Local Agents
Each local agent is responsible for deciding the placement of application components onto available resource nodes within its local network. The action space for a local agent is represented as:
@@ V-C1 Action Space for Local Agents
Where:
  • $a_{ij}=1$ indicates that application component $c_{i}$ is placed on resource node $v_{j}$.
  • $a_{ij}=0$ indicates that $c_{i}$ is not placed on $v_{j}$.
  • $V_{local}$ is the set of resource nodes within the local agent’s neighborhood.
To accommodate dynamic environments and support scalability, a masked action space is employed: a certain percentage of additional resource nodes are initialized but marked as unavailable for current decision-making, reserving them for potential future dynamic resource addition. Resource availability is dynamically updated over time dependent on changing application and environmental factors.
V-C2 Action Space for Global Agent
The global agent focuses on high-level decisions by selecting local agents to execute resource management tasks. Its action space is represented as:
@@ V-C2 Action Space for Global Agent
Where:
  • $a_{i}=1$ indicates that the global agent delegates resource management tasks to the $i$-th local agent.
  • $a_{i}=0$ indicates that the global agent does not delegate tasks to the $i$-th local agent.
  • $N_{local}$ is the total number of local agents in the system.
The proposed strategy offers decentralized management, where local agents handle granular task placement within their neighborhoods while the global agent manages high-level coordination. To meet large-scale requirements, future work will explore local agents collaboration by integrating the action and state spaces of multiple agents, enabling more coordinated and efficient decision-making across the system.
@@ V-D Optimization Techniques
V-D1 GNN Embeddings
GCN or GraphSAGE mentioned in III-B may be used to encode the graph states of the resource network and application components. These embeddings provide a compact representation of the system, aiding decision-making.
V-D2 Multi-Agent DRL Optimization
MARL methods may be explored to optimize application placement and resource allocation in a distributed setting. Algorithms such as A3C (Asynchronous Advantage Actor-Critic and PPO (Proximal Policy Optimization) may be implemented from Stable-Baselines3[26] library to streamline development and evaluation.
V-D3 Customized Experience Replay
A customized experience replay strategy may be utilized during the joint training phase to improve training stability and performance for local agents:
  • Replay Buffer: Stores experiences from the pretraining phase of local agents, including state-action-reward tuples.
  • Stabilization Mechanism: Experience replay helps balance updates between pretrained phase and joint phase, preventing instability and promoting faster convergence.
V-E Training Algorithm
The training process consists of two phases: pretraining the local agents and jointly training the global agent alongside the local agents, as detailed in 1. Below, we highlight the key features of the training algorithm.
Algorithm 1 Training Algorithm for Local and Global Agents with Sequential Placement
0:  Resource graph $G_{R}$, Application graph $G_{A}$
0:  Trained local policy, $\pi_{local_{i}}$ and global policy $\pi_{global_{i}}$
1:  Phase 1: Pretraining Local Agents
2:  for each local agent $i$ do
3:     Extract local resource subgraph $G_{R}^{local}(i)\subseteq G_{R}$
[…]
5:     Apply a sequential placement strategy:
6:     for each application component $c\in G_{A}$ do
7:        Select a resource node $v\in G_{R}^{local}(i)$ using $\pi_{local_{i}}$
8:        Map communication requirements using a shortest path algorithm to determine the optimal communication path
9:     end for
10:     Train local agent $i$ using RL optimization technique to optimize the local reward: $R_{local_{i}}$
11:     Update local policy $\pi_{local_{i}}$
12:  end for
13:  Phase 2: Joint Training of the Global Agent
[…]
15:  Apply a sequential placement strategy:
16:  for each application do
17:     Select a local agent $i$ using $\pi_{global}$
18:     update $O_{global}$ by repeating step 3 to 8 for local agent $i$
19:  end for
20:  Joint training the global agent and local agents using RL optimization technique to optimize the combined reward:$R_{global}=\lambda\cdot R_{non-local}+\sum_{i}\mu_{i}\cdot R_{local_{i}}$
21:  Update global policy $\pi_{global}$ and local policies $\pi_{local_{i}}$
22:  return  Final trained policies $\pi_{local_{i}}$ for all local agents and $\pi_{global}$ for the global agent
  • Sequential Placement Strategy: At each time step, the local agent assigns the current application component to a resource node. Communication requirements such as latency and bandwidth are validated using a shortest path algorithm to ensure placement compliance.
  • Integration of GNN Embeddings: GNN embeddings are employed by both local and global agents to represent their observation spaces, providing a deeper understanding of the resource and application states.
  • MARL Training: Pretrained local agents collaborate with a global agent to jointly optimize application placement and resource allocation in the Cloud-Edge Continuum.
@@ VI-A Simulation Environment
We propose using a simulation framework to verify the approach. A suitable simulator should support:
  • Modeling of the Cloud-Edge Continuum, including resource nodes (e.g., cloud servers, edge devices, IoT nodes) and their communication links.
  • Configurable network topologies to evaluate performance under various scenarios.
  • Support for multi-agent reinforcement learning (MARL) with customizable policies for local and global agents.
  • Scalability to simulate large-scale systems with dynamic task arrivals and resource variability.
We are considering extending an existing simulator described in III-E to create a tailored platform for implementing and testing MARL in distributed environments. This enables fine-grained control over simulation parameters and integration of specific features required for evaluating the proposed approach.
@@ VI-B Evaluation Metrics
The following metrics will be used to assess the performance of the proposed system:
  • SLA Violation Rate: The percentage of application tasks that meet their specified deadlines.
  • Global Resource Utilization: The average utilization of computational and communication resources across all nodes and edges.
  • Completion Time: The total time taken to complete application placement and execution.
  • Training Time and Efficiency: The time required to train local and global agents, as well as the efficiency of the training process.
  • Memory Usage for Training: The memory consumption during the training phase, highlighting the feasibility of scaling to larger systems.
VI-C Comparison with Centralized Approaches
The performance of the proposed decentralized multi-agent framework will be compared against traditional centralized resource management approaches. Key comparison points include: scalability to large-scale environments; adaptability to dynamic changes in network topology and resource availability; overall performance improvements in SLA adherence, resource utilization, and task completion times.
VII Conclusions
We presented our viewpoint on exploiting distributed local and global learning for scalable resource management in Continuum platform, presenting a hierarchical system organization and a training algorithm for the local and global agents. The work to validate our approach via implementation and simulation of a Continuum platform is still ongoing. Building on the findings of the practical evaluation, we plan to extend and validate our approach further through the following steps:
  • Different Network Topologies: Investigate performance across various network configurations, including hierarchical, mesh, and hybrid topologies.
  • Scaling Up: Develop hierarchical coordination and validate system scalability through simulations in larger-scale environments with more nodes, edges, and application components.
  • Improving cost functions: The completion time cost model from section V assumes that application components run sequentially as an upper bound. Literature results about behavioral models of parallel and distributed patterns will be exploited to improve the completion time models.
  • Real-World Testing: Use real world testbeds to validate the practicality and robustness of the proposed system.
  • Dynamic Agent Creation: Explore mechanisms for dynamically creating and adapting agents to accommodate changes in network scale and task demands.
  • Economic Costs Inclusion: Incorporate economic considerations into the optimization framework, comparing cost-effectiveness of centralized and decentralized approaches.

### v2

@@ V-A2 Global Agent
  • Aggregate Observability: Receives abstracted state information from local agents, detailed in Sect.V-B.
  • Joint Training: Trained with local agents to align local and global objectives.
  • Global Reward Considers non-local and local metrics:
@@
    Where $R_{non-local}$ is defined similarly to $R_{local}$ but for system-wide metrics and $\lambda,\mu_{i}$ weight the importance of non-local and local objectives.
@@ V-B Observation Space
Local Agent Observation
Each local agent has partial observability: it only sees the subgraph of $G_{R}$ corresponding to its local neighbourhood plus the full application graph. More formally:
$O_{local}=\{S_{app},S_{res}^{local}\}$ (4)
where the application state visible to all agents is defined as:
@@
and the resource subgraph assigned to that local agent is:
@@
Combined Observation Space for the Global Agent
The global agent maintains a high-level view of the entire system by aggregating local subgraph information:
@@
where $S_{app}$ is the same overall application state as seen by the local agents and $S_{res}$ is the aggregated resource view composed of all local subgraphs.
@@ V-C Action Space
  • Local Agents: Assign application components $c_{i}$ to resource nodes $v_{j}$ within their subgraph:
@@
    Here $a_{ij}=1$ indicates that $c_{i}$ is placed on $v_{j}$.
  • Global Agent: Decides which local agents to activate for resource management tasks:
@@
    where $N_{local}$ is the number of local agents in the system.
The proposed strategy offers decentralized management, where local agents handle granular task placement within their neighborhoods while the global agent manages high-level coordination.
@@ V-D Optimization Techniques
GNN Embeddings: Methods like GraphSAGE compress node and edge information for both local and global agents.
Multi-Agent DRL: Algorithms such as PPO and A3C coordinate decisions decisions across the Cloud-Edge Continuum, leveraging frameworks like Stable-Baselines3
Experience Replay: Stores experiences from pretraining and joint training for more stable and rapid convergence.
V-E Training Overview
The training process consists of two phases: pretraining the local agents and jointly training the global agent alongside the local agents.
Phase 1: The GNN generates embeddings to create $O_{local_{i}}$, and local components are placed sequentially using local policies. The local agent policies are trained MARL.
Phase 2: The GNN generates embeddings by aggregating local information to represent the global state. The global agent then sequentially selects the local agent to execute the sequential placement strategy. The global and pretrained local agents are jointly trained to optimize a combined reward.
@@ VI-A Simulation Environment
A simulation framework like ECLYPSE[9] could be used to verify the approach. Used for distributed computing systems, with an emphasis on edge-cloud infrastructure, it is a highly extensible framework and offers detailed application behavior modeling. We are considering extending it to create a tailored platform for implementing and testing MARL in distributed environments, enabling fine-grained control over simulation parameters.
@@ VI-B Evaluation Metrics
The metrics previously defined in IV-A3, RU, CT and SVR, will be used to assess the performance of the proposed system. Additionally, we will use Training Efficiency, defined as the computational efficiency of the training process including Training Time and Memory Usage, highlighting the feasibility of scaling to larger systems.
VII Conclusion
We presented our viewpoint on exploiting distributed local and global learning for scalable resource management in the Continuum platform, introducing a hierarchical system organization and an overview of the training process for the local and global agents. Building on the findings of the practical evaluation, we plan to validate our approach further through several steps. First, we aim to use real-world testbeds to validate the practicality and robustness of the proposed system. Second, we intend to develop hierarchical coordination and validate the system’s scalability through simulations in larger-scale environments with more nodes, edges, and application components. Third, we plan to incorporate economic considerations into the optimization framework, comparing the cost-effectiveness of centralized and decentralized approaches.
To meet large-scale requirements, future work will explore local agents collaboration by integrating the action and state spaces of multiple agents, enabling more coordinated and efficient decision-making across the system.